OpenStudy (anonymous):

find a closed form of nth partial sums and determined whether the series in converges k=1sigma k tends to infinity (2/(5^(k-1))

OpenStudy (anonymous):

plz help me..and one help me.:(

OpenStudy (anonymous):

any one there..

OpenStudy (anonymous):

firstly closed form of the serie \[\sum_{k=1}^{n} \frac{2}{5^{k-1}}=2 \sum_{k=1}^{n} \frac{1}{5^{k-1}}=2(1+\frac{1}{5}+\frac{1}{5^2}+\frac{1}{5^3}+...+\frac{1}{5^{n-1}})\] thats a geometric serie

OpenStudy (anonymous):

do u know what is the closed form of geometric serie?

OpenStudy (anonymous):

no brother i dont know i saw first time tthis series

OpenStudy (anonymous):

what is the closed form of nth partial sums ? can u explain..?

OpenStudy (anonymous):

well for geometric serie like a+aq+aq^2+aq^3+...aq^(n-1) closed form is \[\frac{a(1-q^n)}{1-q}\]

OpenStudy (anonymous):

closed form of nth partial sums is the partial sum of all the terms from the first and up to the nth which is aq^(n-1)

OpenStudy (anonymous):

in my question what is the cloased form of nth partial sum>?

OpenStudy (anonymous):

ok u have a=1 and q=1/5 then closed form of nth partial sums is \[\frac{1-(\frac{1}{5})^n}{1-\frac{1}{5}}\]

OpenStudy (anonymous):

i forgot the constant coefficient 2

OpenStudy (anonymous):

actually u will have \[S _{n}=2\frac{1-(1/5)^n}{1-1/5}\]

OpenStudy (anonymous):

whether the series is converges..?

OpenStudy (anonymous):

well for that case u have to calculate \[\lim_{n \rightarrow \infty} S _{n}\] We will say that series is convergent if and only if the lim is exist

OpenStudy (anonymous):

now u have \[\lim_{n \rightarrow \infty} 2\frac{1-(1/5)^n}{1-1/5}=\frac{2}{1-1/5}=\frac{10}{4}=2.5\]

OpenStudy (anonymous):

then the series is converges

OpenStudy (anonymous):

(1/5)^ infinity is zero..?

OpenStudy (anonymous):

yes thats right

OpenStudy (anonymous):

brother where is series converges or diverges

OpenStudy (anonymous):

brother in this question firsr four partial sum is 310/125 my ans is right or wrong.?

OpenStudy (anonymous):

well thats right

OpenStudy (anonymous):

brother any direct formula u know tho find series converges or diverges.?

OpenStudy (anonymous):

no... any direct formula

OpenStudy (anonymous):

bhahi jaza ALLAH.. allah give u many happiness thanks

OpenStudy (anonymous):

welcome